# TNPAR: Topological Neural Poisson Auto-Regressive Model for Learning Granger Causal Structure from Event Sequences

### v1

@@ 3.2 Generation Process of the Event Sequences via Topological Neural Poisson Auto-Regressive Model
3.3 Inference Process of the Granger Causal Structure via Amortized Inference
The inference process of TNPAR is depicted using dashed lines in Fig. 2. As shown in the diagram, the Granger causal structure $\mathbf{A}_{0:K}$ is inferred based on a combination of the current occurrence number $O_{t}^{v_{i},n_{j}}$, historical occurrence numbers $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ and the topological network $\mathbf{B}_{0:K}$. Recently, amortized inference, as an advanced method of variational inference, has garnered substantial interest from scholars Zhang et al. (2018); Gershman and Goodman (2014). This approach leverages powerful functions such as neural networks to generate variational distributions conditioned on input data, thereby reducing the computational cost compared to traditional variation inference. Consequently, we designate the Granger causal structure as the latent variable of amortized inference to implement the inference process.
To amortize the inference task for distribution of data $\mathrm{P}(\mathbf{O}^{V,N})$, we optimize an inference model $\mathrm{h}_{\gamma}$ to predict the variational posterior distribution of Granger causal structure $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$. Specifically, given the data $O_{t}^{v_{i},n_{j}}$ and $\mathbf{O}^{V,N}_{t-1:t-\Omega}$, as well as the topological network $\mathbf{B}_{0:K}$, we use a recognition model $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$ to approximate the true posterior $\mathrm{p}_{\theta}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$. Here, $\phi$ and $\theta$ represent the variational and generation parameters, respectively.
To learn the optimal values of parameters $\theta$ and $\phi$, we model each item in the occurrence numbers $\mathbf{O}^{V,N}$ with a Poisson process, as described in Eq. (1). Subsequently, we maximize the ensuing likelihood function for each $O_{t}^{v_{i},n_{j}}$:
@@ 3.3 Inference Process of the Granger Causal Structure via Amortized Inference
where $\mathrm{q}^{A}_{\phi}=\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|\mathrm{h}_{\gamma}(O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega}),\mathbf{B}_{0:K})$, $\mathrm{p}^{A}_{\theta}=\mathrm{p}_{\theta}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$. In Eq. (5), the right-hand side’s (RHS) first term represents the Kullback-Leibler (KL) divergence of the approximate posterior from the true posterior, and the RHS’s second term is the variational evidence lower bound (ELBO) on the log-likelihood of $O_{t}^{v_{i},n_{j}}$. By defining $\mathcal{L}_{e}$ as the ELBO $\mathcal{L}(\theta,\phi;O_{t}^{v_{i},n_{j}}|\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$, and considering that the KL divergence is non-negative, we can deduce from the Eq. (5) that:
@@
Furthermore, the ELBO can be expressed in the following manner:
@@
3.4 Theoretical Analysis
In this subsection, we will theoretically analyze the identifiability of the proposed model. To do so, we begin with the definition, according to work in Tank et al. (2018), of the Granger non-causality of multi-type event sequences as,
@@ 3.4 Theoretical Analysis
Given data $\mathbf{O}^{V}$ during time interval $t\in\{1,\dots,T/\Delta\}$, which is generated according to Eq. (2), we can determine the Granger non-causality of event type $v_{i}$ with respect to event type $v_{j}$ as follows: For all $\mathbf{O}^{v_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|}}_{t-1:t-\Omega}$ and $\mathbf{O}^{v_{i}}_{t-1:t-\Omega}\neq\mathbf{O}^{v_{i}^{{}^{\prime}}}_{t-1:t-\Omega}$, if the following condition holds:
@@
$=$ $\mathrm{g}(\mathbf{O}^{v_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{i}^{{}^{\prime}}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|}}_{t-1:t-\Omega}).$
That is, $O_{t}^{v_{j}}$ is invariant to $O_{t}^{v_{i}}$ with g()normal-$\mathrm{g}(\cdot)$.
@@
Assume that the occurrence numbers of event sequences are generated by Eq. (4). We can determine the Granger non-causality of event type $v_{i}$ with respect to event type $v_{j}$ as follows: For all $\mathbf{O}^{v_{1},n_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|},n_{|N|}}_{t-1:t-\Omega}$ and $\mathbf{O}^{v_{i},n_{k}}_{t-1:t-\Omega}\neq\mathbf{O}^{v_{i}^{{}^{\prime}},n_{k}}_{t-1:t-\Omega},k\in\{1,...,|N|\}$, if the following condition holds in all nodes under topological network $\mathcal{G}_{N}$:
@@
$=$ $\mathrm{g}(\mathbf{O}^{v_{1},n_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{i}^{{}^{\prime}},n_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{i}^{{}^{\prime}},n_{|N|}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|},n_{|N|}}_{t-1:t-\Omega}).$
That is, $O_{t}^{v_{j}}$ is invariant to $O_{t}^{v_{i}}$ with g()normal-$\mathrm{g}(\cdot)$ in all nodes under topological network $\mathcal{G}_{N}$.
@@
Given the Granger causal structure $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$, the topological network $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$ and the max geodesic distance $K$, along with the assumption that the data generation process adheres to Eq. (4), we deduce that $v_{i}\rightarrow v_{j}\notin\mathbf{E}_{V}$, if and only if $A_{k}^{v_{i},v_{j}}=0$ for every $k\in\{0,...,K\}$.
The comprehensive proof of Proposition 1 can be found in Appendix B. This proposition inspires a methodology for TNPAR to identify Granger causality, which involves evaluating whether the elements of $\mathbf{A}_{0:K}$ are zero.
4 Algorithm and Implementation
4.1 Inference Process of TNPAR
In this subsection, we employ an encoder to implement the above inference process of TNPAR. Here, the present data $O_{t}^{v_{i},n_{j}}$ and the historical data $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ are used jointly as inputs to the encoder, with $\mathbf{B}_{0:K}$ serving as a mask. The output is the variational posterior of the Granger causal structure, represented as $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$.
To elaborate, as described in Eq. (4), each element within $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ is initially filtered in accordance with both $\mathbf{B}_{0:K}$ and $n_{j}$. The resulting values are then combined with $O_{t}^{v_{i},n_{j}}$ and converted into a vector, where each item denotes the representation of the occurrence number at a geodesic distance of $k$ for the corresponding node. Following this, the encoder applies a multilayer perception (MLP) Hastie et al. (2009); Cybenko (1989) to the input, which facilitates the propagation of input information across multiple layers and nonlinear activation functions. Ultimately, the encoder generates the posterior distribution of Granger causal structure, denoted as $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|\mathrm{h}_{\gamma}(O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega}),\mathbf{B}_{0:K})$.
In a manner similar to variational inference for categorical latent variables, we assume that the posterior distribution of each causal edge in the Granger causal structure follows a Bernoulli distribution. More specifically, the encoder produces a latent vector $\mathbf{Z}\in[0,1]^{|\mathbf{V}|\times|\mathbf{V}|\times K}$, where $z_{i,j,k}$, the $(i\times j\times k)$-th element of $\mathbf{Z}$, represents the Bernoulli distribution parameter for corresponding edge. Consequently, $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|\mathrm{h}_{\gamma}(O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega}),\mathbf{B}_{0:K})$ can be expressed as:
$\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|\mathrm{h}_{\gamma}(O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega}),\mathbf{B}_{0:K})$ (8)
@@ 4.1 Inference Process of TNPAR
Furthermore, for backpropagation through the samples of the discrete distribution $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|\mathrm{h}_{\gamma}(O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega}),\mathbf{B}_{0:K})$ while training TNPAR, we employ the Gumbel-Softmax trick Jang et al. (2017) to generate a differentiable sample as:
$\hat{A}_{k}^{v_{i},v_{j}}=\frac{\exp(\log(\mathrm{q}_{\phi}(A_{k}^{v_{i},v_{j}}=1|z_{i,j,k})+\varepsilon_{1})/\tau)}{\sum_{e=0}^{1}\exp(\log(\mathrm{q}_{\phi}(A_{k}^{v_{i},v_{j}}=e|z_{i,j,k})+\varepsilon_{e})/\tau)}$,
where each $\varepsilon_{*}$ is an i.i.d. sample drawn from the standard Gumbel distribution, while $\tau$ is the temperature parameter controlling the randomness of $\hat{A}_{k}^{v_{i},v_{j}}$.
4.2 Generation Process of TNPAR
Next, we propose a decoder to implement the generation process of $\mathbf{O}_{t}^{V,N}$. In this part, $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ is filtered according to both the estimated $\hat{\mathbf{A}}_{0:K}$ and the prior topological network $\mathbf{B}_{0:K}$, as specified in Eq. (4). This filtered result is then used as input to the decoder. Notably, for neural network training, we treat $\hat{\mathbf{A}}_{0:K}$ as a weight matrix during training, and as a mask during testing. This can be accomplished using soft/hard Gumbel-Softmax sampling. Similar to the encoder, a Multilayer Perceptron (MLP) is utilized in the decoder.
Aligning with numerous works on counting processes that posit the distribution of occurrence numbers $O_{t}^{v_{i},n_{j}}$ in each time interval as stationary Stoyan et al. (2013); Babu and Feigelson (1996); Othmer et al. (1988), we model each $\mathrm{p}_{\theta}(O_{t}^{v_{i},n_{j}}|\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{A}_{0:K},\mathbf{B}_{0:K})$ using a Poisson process as denoted in Eq. (1). In this context, the conditional distribution of $O_{t}^{v_{i},n_{j}}$ can be formalized as:
@@ 4.2 Generation Process of TNPAR
4.3 Optimization of TNPAR
Based on the above analysis, we adopt the variational ELBO of variational inference, defined in Eq. (7), as the objective function to estimate parameters $\phi$ and $\theta$ of TNPAR and infer the posterior of latent Granger causal structure. In this context, the RHS’s first term of Eq. (7) represents the reconstruction loss, while the RHS’s second term signifies the Kullback-Leibler loss.
In numerous real-world applications, there is often no causal loop in the causal structure $\mathcal{G}_{V}$ during the generation process of event sequences. For instance, based on experts’ prior knowledge, the causal structure in the aforementioned operation and maintenance scenario of a mobile network is a DAG (Directed Acyclic Graph). With this in mind, we introduce an acyclic constraint proposed by Yu et al. (2019), denoted as:
@@ 4.3 Optimization of TNPAR
where $\mathbf{G}$ is a matrix in which $G_{i,j}=0$ signifies that event type $v_{i}$ has no Granger causality with $v_{j}$. Conversely, $G_{i,j}>0$. The function $\mathrm{tr}(\cdot)$ calculates the trace of an input matrix. A causal structure $\mathbf{G}$ is acyclic if and only if $\mathrm{h}(\mathbf{G})=0$; Otherwise, $\mathrm{h}(\mathbf{G})>0$. In this study, we set $G_{i,j}=\sum_{k=0}^{K}A_{k}^{v_{i},v_{j}}$ and populate the diagonal of $\mathbf{G}$ with 00, which implies that self-excitation is permitted in an acyclic graph. Subsequently, we design an acyclic regular term $\mathcal{L}_{c}$ as:
@@
Without loss of generality, the causal structure tends to be sparse in most real-world applications. Thus, we employ a $\ell_{1}$-norm sparsity constraint to control the model’s sparsity as follows:
@@
where $\kappa$ is a small positive constant. Therefore, the training procedure of TNPAR boils down to the following constrained optimization:
@@
By leveraging the Lagrangian multiplier method, we define the total loss function $\mathcal{L}_{total}$ of TNPAR as:
@@
where both $\lambda_{c}$ and $\lambda_{s}$ denote the regularization hyperparameters.
In summary, our proposed model is trained on the occurrence numbers of event sequences using the following objective:
@@
For model training, we can adopt Adam Kingma and Ba (2015) stochastic optimization.
5 Experiment
In this section, we conduct experiments, applying the proposed TNPAR method and the various baseline methods to both simulated data and real metropolitan cellular network alarm data. For all methods, we employ five different random seeds and present the results in graphical format, with error bars included. Our evaluation metrics for the experiments include Precision, Recall, F1 score Powers (2011), Structural Hamming Distance (SHD), and Structural Intervention Distance (SID) Peters and Bühlmann (2015). Specifically, Precision refers to the fraction of predicted edges that exist among the true edges. Recall is the fraction of true edges that have been successfully predicted. F1 score is the weighted harmonic mean of both Precision and Recall, and is calculated as $F1=\frac{2\times Precision\times Recall}{Precision+Recall}$. SHD represents the number of edge insertions, deletions, or flips needed to transform a graph into another graph. SID is a measure that quantifies the closeness between two DAGs based on their corresponding causal inference statements.
5.1 Dataset
5.1.1 Simulated Data
The simulated data is generated as follows: a) a Granger causal structure $\mathcal{G}_{V}$ and a topological network $\mathcal{G}_{N}$ are randomly generated; b) root event records are generated via the Poisson process using a base intensity parameter $\mu$ in the Hawkes process. These root event records are spontaneously generated by the system; c) based on the root event records, propagated event records are discretely generated according to both the time interval $\Delta$ and the excitation intensity $\alpha$. Here, $\alpha$ represents the event excitation intensity in the Hawkes process. Given that the event sequences can be quite sparse in real-world data, we offer a time interval parameter $\Delta$ to the generation process, dividing the time domain $[0,T]$ to $[0,T/\Delta]$. In this context, the event records can be summarized within the same timestamp. Note that the $\Delta\geq 0$. If $\Delta=0$, it implies the use of the original event sequences.
@@ 5.1.1 Simulated Data
5.1.2 Metropolitan Cellular Network Alarm Data
The Metropolitan Cellular Network alarm event data, which was collected in a business scenario by a multinational communications company, is available in the PCIC 2021 competition. This data comprises a series of alarm records generated within a week according to both a topological network $\mathcal{G}_{N}$ and a causal structure $\mathcal{G}_{V}$. Specifically, the topological network $\mathcal{G}_{N}$ includes $3087$ network elements. The alarm records encompass $18$ alarm event types, totaling $228,030$ alarm event records. It is noteworthy that, due to the characteristics of the equipment and as confirmed by experts, the collection times of alarm events exhibit certain time intervals.
@@ 5.1.2 Metropolitan Cellular Network Alarm Data
5.2 Baselines and Model Variants
The following Granger causal discovery methods are chosen as baseline comparisons in our experiments: PCMCI Runge (2020), MLE-SGL Xu et al. (2016), ADM4 Zhou et al. (2013), and THP Cai et al. (2022). In order to assess the components of our proposed method, we also introduce two variants of our method, namely TNPAR_NT and TNPAR_NC. Specifically, TNPAR_NT is a variant that does not incorporate topological information, while TNPAR_NC is a version that excludes both the acyclicity and sparsity regularized terms from the loss function. Detailed information about the baseline methods can be found in Appendix C.
5.3 Results on Simulated Data

### v2

@@ Introduction
Inference Process of the Granger Causal Structure via Amortized Inference
The inference process of TNPAR is depicted using dashed lines in Fig. 2. As shown in the diagram, the causal matrices $\mathbf{A}_{0:K}$ are inferred based on a combination of the current occurrence number $O_{t}^{v_{i},n_{j}}$, historical occurrence numbers $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ and the topological network $\mathbf{B}_{0:K}$. Note that the causal relationship is invariant across the varying samples. This nature inspires us to implement the inference process with amortized inference (Zhang et al. 2018; Gershman and Goodman 2014), in which a shared function is employed to predict the varying posterior distribution of each sample. Consequently, we consider the generation process function, defined in Eq. (4), as the shared function and designate the causal matrices as the function’s parameter to implement the amortized inference process.
In more detail, to amortize the inference of the shared function, we optimize an inference model to predict the distribution of causal matrices. Given the data $O_{t}^{v_{i},n_{j}}$ and $\mathbf{O}^{V,N}_{t-1:t-\Omega}$, as well as the topological network $\mathbf{B}_{0:K}$, we use $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$ to approximate the true distribution $\mathrm{p}_{\theta}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$. Here, $\phi$ and $\theta$ represent the variational and generation parameters, respectively. The value of $\theta$ and $\phi$ can be learned by maximizing the ensuing likelihood function for each $O_{t}^{v_{i},n_{j}}$:
@@
where $\mathrm{q}^{A}_{\phi}=\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$, $\mathrm{p}^{A}_{\theta}=\mathrm{p}_{\theta}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$. In Eq. (5), the right-hand side’s (RHS) first term represents the Kullback-Leibler (KL) divergence of the approximate posterior from the true posterior, and the RHS’s second term is the variational evidence lower bound (ELBO) on the log-likelihood of $O_{t}^{v_{i},n_{j}}$. By defining the ELBO as $\mathcal{L}_{e}$ and considering the KL divergence is non-negative, we can deduce from the Eq. (5) that:
@@
The ELBO can be further expressed as:
@@
Theoretical Analysis
In this subsection, we will theoretically analyze the identifiability of the proposed model. To do so, according to the definition in (Tank et al. 2018), we begin with the definition of the Granger non-causality of multi-type event sequences as follows:
@@
Given data $\mathbf{O}^{V}$ during time interval $t\in\{1,\dots,\lceil T/\Delta\rceil\}$, which is generated according to Eq. (2). For the original sample values of historical variables $\mathbf{O}^{v_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|}}_{t-1:t-\Omega}$, let $\mathbf{\hat{O}}^{v_{i}}_{t-1:t-\Omega}$ represent the different sample values of the historical variables $\mathbf{O}^{v_{i}}_{t-1:t-\Omega}$, (i.e., $\mathbf{O}^{v_{i}}_{t-1:t-\Omega}\neq\mathbf{\hat{O}}^{v_{i}}_{t-1:t-\Omega}$). We can determine the Granger non-causality of event type $v_{i}$ with respect to event type $v_{j}$, if the following condition holds for all $t$:
@@
$=$ $\mathrm{g}(\mathbf{O}^{v_{1}}_{t-1:t-\Omega},...,\mathbf{\hat{O}}^{v_{i}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|}}_{t-1:t-\Omega}).$
That is, $O_{t}^{v_{j}}$ is invariant to $\mathbf{O}^{v_{i}}_{t-1:t-\Omega}$ with g()normal-$\mathrm{g}(\cdot)$.
@@
Assume that the occurrence numbers of event sequences are generated by Eq. (4). For the original sample values of historical variables $\mathbf{O}^{v_{1},n_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|},n_{|N|}}_{t-1:t-\Omega}$, let $\mathbf{\hat{O}}^{v_{i},n_{k}}_{t-1:t-\Omega}$ represent the different sample values of the historical variables $\mathbf{O}^{v_{i},n_{k}}_{t-1:t-\Omega}$, (i.e., $\mathbf{\hat{O}}^{v_{i},n_{k}}_{t-1:t-\Omega}\neq\mathbf{O}^{v_{i},n_{k}}_{t-1:t-\Omega}$). We can determine the Granger non-causality of event type $v_{i}$ with respect to event type $v_{j}$ under topological network $\mathcal{G}_{N}$, if the following condition holds for all $t$:
@@
$=$ $\mathrm{g}(\mathbf{O}^{v_{1},n_{1}}_{t-1:t-\Omega},...,\mathbf{\hat{O}}^{v_{i},n_{1}}_{t-1:t-\Omega},...,\mathbf{\hat{O}}^{v_{i},n_{|N|}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|},n_{|N|}}_{t-1:t-\Omega}).$
That is, $O_{t}^{v_{j},n_{k}}$ is invariant to $\mathbf{O}^{v_{i},n_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{i},n_{|N|}}_{t-1:t-\Omega}$ with g()normal-$\mathrm{g}(\cdot)$ across all nodes in topological network $\mathcal{G}_{N}$.
@@
Given the Granger causal structure $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$, the topological network $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$ and the max geodesic distance $K$, along with the assumption that the data generation process adheres to Eq. (4), we can deduce that $v_{i}\rightarrow v_{j}\notin\mathbf{E}_{V}$, if and only if $A_{k}^{v_{i},v_{j}}=0$ for every $k\in\{0,...,K\}$.
The proof of Proposition 2 is provided in Appendix B. This proposition inspires a methodology for TNPAR to identify Granger causality, which involves evaluating whether the elements of $\mathbf{A}_{0:K}$ are zero.
Algorithm and Implementation
Inference Process of TNPAR
In this subsection, we employ an encoder to implement the above inference process of TNPAR. Here, the $O_{t}^{v_{i},n_{j}}$ and $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ are used jointly as inputs to the encoder, with $\mathbf{B}_{0:K}$ serving as a mask. The output is the variational posterior of the Granger causal structure, represented as $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$.
To elaborate, each element within $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ is initially filtered in accordance with both $\mathbf{B}_{0:K}$ and $n_{j}$. The resulting values are then combined with $O_{t}^{v_{i},n_{j}}$ and converted into a vector, where each item denotes the representation of the occurrence numbers at a geodesic distance of $k$ for $n_{j}$. Following this, the encoder applies a multilayer perception (MLP) (Hastie et al. 2009) to the input, which facilitates the propagation of input information across multiple layers and nonlinear activation functions. Ultimately, the encoder generates the posterior distribution of Granger causal structure $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$.
In a manner similar to variational inference for categorical latent variables, we assume that the posterior distribution of each causal edge in the Granger causal structure follows a Bernoulli distribution. More specifically, the encoder produces a latent vector $\mathbf{Z}\in[0,1]^{|\mathbf{V}|\times|\mathbf{V}|\times K}$, where $z_{i,j,k}$, the $(i\times j\times k)$-th element of $\mathbf{Z}$, represents the parameter of Bernoulli distribution for corresponding edge. Consequently, $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$ can be expressed as:
$\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$ (8)
@@
Furthermore, for backpropagation through the discrete distribution $\mathrm{q}_{\phi}(\mathbf{A}_{0:K}|O_{t}^{v_{i},n_{j}},\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$, we employ the Gumbel-Softmax trick (Jang, Gu, and Poole 2017) to generate a differentiable sample as:
$\hat{A}_{k}^{v_{i},v_{j}}\!=\!\frac{\exp(\log(\mathrm{q}_{\phi}(A_{k}^{v_{i},v_{j}}\!=\!1|z_{i,j,k})+\varepsilon_{1})/\tau)}{\sum_{e=0}^{1}\exp(\log(\mathrm{q}_{\phi}(A_{k}^{v_{i},v_{j}}\!=\!e|z_{i,j,k})+\varepsilon_{e})/\tau)},$ (9)
where each $\varepsilon_{*}$ is an i.i.d. sample drawn from a standard Gumbel distribution, while $\tau$ is the temperature parameter controlling the randomness of $\hat{A}_{k}^{v_{i},v_{j}}$.
Generation Process of TNPAR
Next, we propose a decoder to implement the generation process of $\mathbf{O}_{t}^{V,N}$. In this part, $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ is filtered according to both $\hat{\mathbf{A}}_{0:K}$ and $\mathbf{B}_{0:K}$, as specified in Eq. (4). This filtered result is then used as input to the decoder. Notably, for neural network training, we treat $\hat{\mathbf{A}}_{0:K}$ as a weight matrix during training, and as a mask during testing. This can be accomplished using soft/hard Gumbel-Softmax sampling. Similar to the encoder, an MLP is used in the decoder.
Aligning with numerous works on counting processes that assume the distribution of occurrence numbers $O_{t}^{v_{i},n_{j}}$ in each time interval is stationary (Stoyan et al. 2013; Babu and Feigelson 1996), we model each $\mathrm{p}_{\theta}(O_{t}^{v_{i},n_{j}}|\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{A}_{0:K},\mathbf{B}_{0:K})$ using a Poisson process as denoted in Eq. (1). In this context, the conditional distribution of $O_{t}^{v_{i},n_{j}}$ can be formalized as:
@@
Optimization of TNPAR
Based on the above analysis, we adopt the ELBO $\mathcal{L}_{e}$, defined in Eq. (7), as the objective function to estimate $\phi$ and $\theta$ and infer the Granger causal structure. To adapt to real-world scenarios, we further introduce acyclic constraint and sparsity constraint into the objective function.
In many real-world applications, the causal structure $\mathcal{G}_{V}$ is acyclic. For instance, based on experts’ prior knowledge, the causal structure in the aforementioned operation and maintenance scenario of a mobile network is a DAG (Directed Acyclic Graph). With this in mind, we introduce an acyclic constraint proposed by (Yu et al. 2019), denoted as:
@@
where $\mathbf{G}$ is a matrix in which $G_{i,j}=0$ signifies that $v_{i}$ has no Granger causality with $v_{j}$. Conversely, $G_{i,j}>0$. The function $\mathrm{tr}(\cdot)$ calculates the trace of an input matrix. Then, a causal structure $\mathbf{G}$ is acyclic if and only if $\mathrm{h}(\mathbf{G})=0$; Otherwise, $\mathrm{h}(\mathbf{G})>0$. In this study, we set $G_{i,j}=\sum_{k=0}^{K}A_{k}^{v_{i},v_{j}}$ and populate the diagonal of $\mathbf{G}$ with 00, which implies that self-excitation is permitted in an acyclic graph. Subsequently, we design an acyclic regulared term $\mathcal{L}_{c}$ as:
@@
Because the causal structure tends to be sparse in most real-world applications, we employ a $\ell_{1}$-norm sparsity constraint as follows:
@@
where $\kappa$ is a small positive constant. Therefore, the training procedure boils down to the following optimization:
@@
By leveraging the Lagrangian multiplier method, we define the total loss function $\mathcal{L}_{total}$ as:
@@
where $\lambda_{c}$ and $\lambda_{s}$ denote the regularized hyperparameters.
In summary, our proposed model is trained using the following objective:
@@
For model training, we can adopt Adam (Kingma and Ba 2015) stochastic optimization.
Experiment
In this section, we apply the proposed TNPAR method and the baselines to both simulated data and metropolitan cellular network alarm data. For all methods, we employ five different random seeds and present the results in graphical format, with error bars included. Our evaluation metrics for the experiments include Precision, Recall, F1 score (Powers 2011), Structural Hamming Distance (SHD), and Structural Intervention Distance (SID) (Peters and Bühlmann 2015). Specifically, Precision refers to the fraction of predicted edges that exist among the true edges. Recall is the fraction of true edges that have been successfully predicted. F1 score is the weighted harmonic mean of both Precision and Recall, and is calculated as $F1=\frac{2\times Precision\times Recall}{Precision+Recall}$. SHD represents the number of edge insertions, deletions, or flips needed to transform a graph into another graph. SID is a measure that quantifies the closeness between two DAGs based on their corresponding causal inference statements.
Dataset
Simulated Data
The simulated data is generated as follows: a) a Granger causal structure $\mathcal{G}_{V}$ and a topological network $\mathcal{G}_{N}$ are randomly generated; b) root event records are generated via the Poisson process using a base intensity parameter $\mu$ in the Hawkes process. These root event records are spontaneously generated by the system; c) based on the root event records, propagated event records are discretely generated according to both the time interval $\Delta$ and the excitation intensity $\alpha$. Here, $\alpha$ represents the event excitation intensity in the Hawkes process. Given that the event sequences can be quite sparse in real-world data, we offer a time interval parameter $\Delta$ to the generation process, dividing the time domain $[0,T]$ to small intervals with indexes as $\{1,\ldots,\lceil T/\Delta\rceil\}$. Then, the event records can be summarized within the same timestamp. Note that the $\Delta\geq 0$. If $\Delta=0$, it implies the use of the original event sequences.
@@
Metropolitan Cellular Network Alarm Data
This real-world dataset is collected in a business scenario by a multinational communications company and is available in the PCIC 2021 competition. The data comprises a series of alarm records generated within a week according to both a topological network $\mathcal{G}_{N}$ and a causal structure $\mathcal{G}_{V}$. Specifically, the topological network $\mathcal{G}_{N}$ includes $3087$ network elements. The alarm records encompass $18$ alarm event types, totaling $228,030$ alarm event records. It is noteworthy that, due to the characteristics of the equipment and as confirmed by experts, the collected timestamps of alarm events exhibit certain time intervals.
@@
Baselines and Model Variants
The following causal discovery methods are chosen as baseline comparisons in our experiments: PCMCI (Runge 2020), ACD (Löwe et al. 2022), MLE_SGL (Xu, Farajtabar, and Zha 2016), ADM4 (Zhou, Zha, and Song 2013), and THP (Cai et al. 2022). In order to assess the components of our proposed method, we also introduce three variants of our method, namely TNPAR_NT, TNPAR_NC, and TNPAR_MG. Specifically, TNPAR_NT does not incorporate topological information and treats each event sequence independently. TNPAR_MG also does not introduce the topological network and takes the event sequences in all nodes as one single sequence. TNPAR_NC is a variant that excludes both acyclicity and sparsity regularization from the loss function. Detailed information about the baselines and variants can be found in Appendix C.
Results on Simulated Data